please can you help me solve this equation :x^2-2x+2=y and 4x-7=y to find x and y thank you very very much .
Answer by ankor@dixie-net.com(22740)

solve this equation :x^2-2x+2=y and 4x-7=y to find x and y
:
Since they both = y, you can write it:
x^2 - 2x + 2 = 4x - 7
:
x^2 - 2x - 4x + 2 + 7 = 0
x^2 - 6x + 9 = 0
:
Factor this to:
(x-3)(x-3) = 0
x = + 3
:
Substitute 3 for x to find y:
y = 3^2 - 2(3) + 2
y = 9 - 6 + 2
y = 5
:
Do the same in y = 4x - 7
y = 4(3) - 7
y = 12 - 7
y = 5 also